OpenStudy (anonymous):

Find the limit of the function algebraically. lim x-->0 x^2 + 3/x^4

OpenStudy (anonymous):

@gurrilabill

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

are there choices

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{x^2+3}{x^4}=\lim_{x\to\infty}\left(\frac{x^2}{x^4}+\frac{3}{x^4}\right)=\lim_{x\to\infty}\left(\frac{1}{x^2}+\frac{3}{x^4}\right)\] Both \(\dfrac{1}{x^2}\) and \(\dfrac{3}{x^4}\) approach \(\infty\) as \(x\to0\) from either direction.